OpenStudy (anonymous):

Which arrangement shows 5 2/5, 5.36, 5 3/5, and 5 3/6 in order from least to greatest?

OpenStudy (triciaal):

Use decimal or fractions for all then compare.

OpenStudy (triciaal):

Do you know how to convert the fraction to decimal?

OpenStudy (anonymous):

the question is in order from least to greatest? im not trying to convert

OpenStudy (anonymous):

@raven7266

OpenStudy (anonymous):

@hhelpplzzzz @horsegal244 @jim_thompson5910

OpenStudy (triciaal):

But you have example 2/5 and 0.36. 2/5 = 0.40 when converted so now compare 36 and 40 to see which is more

OpenStudy (anonymous):

what no 27/5 and 5.36

OpenStudy (raven7266):

least to greatest a. 5.36, (27/5), (33/6), 5(3/5)

OpenStudy (anonymous):

thanks

OpenStudy (raven7266):

is this all your question for this topic?

OpenStudy (triciaal):

My "draw" is not working. My response is limited. Can you picture the number line? All the numbers you have are between 5 and 6 so basically ignore the 5 You have points you have to line up but you need to use the same "markers" Each must have the same value. Here each is 1/100. 2/5 =40/100. So 5.40. Another point is 5.36.

OpenStudy (anonymous):

its not a number line and its closed

OpenStudy (anonymous):

plus i posted some pictures no number lines its not a number line question

OpenStudy (triciaal):

You don't appreciate learning but that's ok. I was showing how you would get the answer. The goal is to know how to get the correct answer.

OpenStudy (triciaal):

Basically yes it is just the number line. Put in order the numbers from least to greatest how they appear on the number line.

OpenStudy (triciaal):

I didn't look at the pictures because those would just be answer choices.

OpenStudy (anonymous):

i do appreciate learning